OpenStudy (jfraser):

The energy and motion of a sample of particles is directly related to their temperature. If you were to cut the temperature of a substance in half, you are basically cutting the eenrgy in half, and cutting the average molecular speed in half, also. What end result does that have on the motion of the molecules and the collisions with the walls?

OpenStudy (anonymous):

the same molecular motion and collisions between particles and the walls of the container?

OpenStudy (anonymous):

is my answer correct ?

OpenStudy (jfraser):

no. If you cut the temp in half, you cut the energy in half, so you cut the motion in half, and the collisions in half.

OpenStudy (anonymous):

The relationship between volume and temperature of a gas, when pressure and moles of a gas are held constant, is: V*T = k. this is true?

OpenStudy (jfraser):

no. When temp increases, volume must also increase. If V*T = k, and they both increase, then k will also increase. The proper order is:\[\frac{V}{T} = k\]

OpenStudy (anonymous):

The relationship between pressure and temperature, when volume and moles of a gas are held constant, is: P/T = k. We could say then, that: If the temperature is doubled, the pressure must remain the same. am i correct?

OpenStudy (anonymous):

these are the choices: If the temperature is doubled, the pressure must be halved. If the temperature is halved, the pressure must also be halved. If the temperature is halved, the pressure must double. If the temperature is doubled, the pressure must remain the same.

OpenStudy (jfraser):

no. If k must be kept constant, and you double T, what must also happen to P?

OpenStudy (anonymous):

If the temperature is doubled, the pressure must be halved.

OpenStudy (jfraser):

no. If temp is doubled, pressure must also double.\[\frac{T}{P} = k = \frac{2*T}{?*P}\] What is the only # that fits in that spot to keep that statement true?

OpenStudy (anonymous):

If the temperature is halved, the pressure must also be halved.

OpenStudy (jfraser):

correct.